OpenStudy (anonymous):

y=x^2-2x-1 y=x+3 determine the common solution of these equations.

OpenStudy (anonymous):

4

OpenStudy (anonymous):

How do you figure that out? @mz_samosino

OpenStudy (anonymous):

x^2-2x-1 = x + 3 x^2 -2x -x 1 -3 =0 factors: (x-4) (x+1) 4,-1 there are two but if you substitute them in both equation the only solution is 4

OpenStudy (anonymous):

Could you please solve the solution for this question? @mz_samosino

OpenStudy (anonymous):

subsitute the value of your y (2nd equation) in the y of 1st equation

OpenStudy (anonymous):

then cross multiply